# Nuanced Safety in Generative AI: How Demographics Shape Responsiveness to Severity

### v1

@@ 4.1 Metrics for the two properties
[…]
(a) Demographic groups of crowd raters by ethnicity
(b) Demographic groups of crowd raters by age
(c) Demographic groups of crowd raters by gender
Figure 2: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for different demographic groups when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender.
We define the following metrics to quantify the strength of the core inequalities for the two properties: Monotonic Precision Area and Weighted Recall Area.
@@
Monotonic Precision Area for stochastic ordering. We note that the probability $P(U=1|S)$ is equivalent to $Precision(S)$, i.e., the precision when classifying items with score $=S$ as unsafe and items with score $\neq S$ as safe. Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$. If we take the area under the curve defined by $Y_{so}(s)=\sum_{i=0}^{s-1}\{Precision(s)-Max_{j=0}^{i}Precision(j)$} at $s=0,1,2,3,\dots,K$, then a high value would mean consistent increases in $P(U=1|S=s)$ without violations of monotonicity. We normalize the area by the maximum possible area, which is given by $\frac{K}{2}*(\frac{K}{2}+1)$ if $K$ is even and $(\frac{K+1}{2})^{2}$ if $K$ is odd. Intuitively, the maximum area is achieved when precision is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold for distinguishing between unsafe ($U=1$) and safe ($U=0$). When the area is below 0, indicating stochastic ordering worse than random guessing, we cap the area to 0 to ensure a non-negative metric. Additionally, when computing $Y_{s}o(s)$, we ignore all the scores $<s$ that are not used by the rater and have undefined precision. Similarly, if $s$ is a score not used by the rater, we take $Y_{s}o(s)=0$.
@@
Weighted Recall Area for discrimination. We note that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe. If we take the area under the curve defined by $Y_{d}(s)=P(S<s|U=0)*Recall(s)$ at $s=0,1,2,3,\dots,K$, then a high value means high $P(S<s|U=0)$ and $P(S=s|U=1)$. This in turn implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property. $Y_{d}(s)$ is essentially the recall of unsafe examples at score $s$, i.e, $Recall(s)$, weighted by the proportion of safe samples correctly discriminated by a score $S<s$. It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $=s$ to unsafe items while correctly assigning scores $S<s$ to safe items.
Since the two metrics are based on rates, we take their harmonic mean to combine them into one metric. Harmonic mean ensures that low performance on any one metric is strongly penalized. When both metrics are high for a rater, it means the rater is responsive to varying levels of severity. On the other hand, a low monotonic precision area with a high weighted recall area suggests that while the rater may be good at coarsely separating items that are of very distinct severities, they are not good at granularly responding to varying levels of severities. A low monotonic precision area may also result from biases in scale usage that causes the rater to not use the full range of the scale, e.g., central tendency bias or extreme responses bias.
@@ 4.2 Contrasting with traditional metrics
There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity. For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve, or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether the relationship between non-expert ratings scores and the reference scores is monotonic. We highlight some limitations of these approaches below:
@@
  • Lack of focus on behaviour at scores: correlation metrics capture general monotonic relationships between scores and underlying severity, if the latter were accessible, but they do not penalize raters who assign higher scores without meaningful increases in corresponding severity at the scores. Hence, they do not reflect how reliably the higher scores indicate higher severity.
  • Fragility to insignificant variations: traditional metrics focus solely on the ordering of scores relative to severity, without taking into account the magnitude of differences between them. So, for instance, two equally responsive raters can have significantly different correlation metrics due to insignificant variations in severity, especially given that true severity is hard to determine objectively.
The way we define and quantify responsiveness addresses these limitations, while offering simplicity, interpretability, and the ability to get insights into how raters utilize different scores on the Likert scale. Nevertheless, as defined in our data model, different types of safety violations might exist, some inherently more or less likely to be labeled unsafe regardless of perceived severity. While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity. In appendix A, we run simulations to show how our proposed metrics and traditional metrics evaluate different scoring patterns.
5 Responsiveness Evaluation of Groups
To demonstrate the utility of our proposed analysis framework with the responsiveness metrics, we apply it to an existing pluralistic dataset of AI safety ratings.
@@ 5.1 Dataset description
We consider the dataset in Rastogi et al. (2024) which contains annotations from people evaluating the safety of generative AI. Concretely, in this dataset, the crowd raters and expert raters assess the safety of a set of prompt-image pairs, where the crowd raters provide a Likert-scale rating from 0 to 4 (where 0 is not harmful and 4 is completely harmful), and expert raters provide binary rating of 0 (safe) or 1 (unsafe) for each prompt-image pair. The crowd raters are recruited based on their demographics. We identify three demographic axes - gender, ethnicity and age. The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group. The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e. their ethnicity, age, and gender. Following Rastogi et al. (2024), we distinguish top-level demographic groups, e.g. East-Asian, from trisectional demographic groups, e.g. Black–GenZ–Man.
The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair. To obtain binary labels $U$ per prompt-image pair, we replicate each prompt-image pair for every expert. For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote (i.e. the mode of scores from all the individual raters belonging to that group) that we refer to as the plurality score - an integer from 0 to 4. In case of ties, we take the most unsafe score to be the plurality score. The distribution of scores provided by crowd raters in the dataset are shown in figure 1. Group cohesion metrics for each of the top-level groups can be found in Appendix tables 1 and 2.
Figure 2 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for different demographic groups of crowd raters when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender
@@ 5.2 Results by trisectional demographic groups
We now evaluate and compare the responsiveness of different demographic groups of crowd raters, both at the trisectional demographic level (e.g., Latinx–GenZ–Man) as well as the top demographic level (e.g., Latinx, Man, etc.).
Figure 3 shows monotonic precision area, weighted recall area, their harmonic mean (hm), Kendall’s $\tau$ for trisectional demographic crowd groups when binary labels $U$ are obtained from expert raters as described in section 4.
@@
Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), and Kendall’s $\tau$ for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Figure 4: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$ for crowd rater population on three violation types when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
@@
We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest monotonic precision area, and consequently, the lowest harmonic means as well. This indicates that higher scores from these trisections correspond the least to higher reference severity as captured by the expert raters. On the other hand though, we note that both the trisections still achieve weighted recall areas comparable to others. This further suggests that while they concur with others on discriminating between distinct levels of severity, they do not respond to the granular severity of violations the same way.
Figure 2(a) further validates the same as we note that the East-Asian top-level demographic group does not exhibit consistent gains in precision when plurality score goes from 1 to 3. Additionally, we see that Kendall’s $\tau$ shows a similar pattern as weighted recall area given that both capture aspects of concordance. But traditional metrics like Kendall’s $\tau$ do not reflect well the ability to stochastically order.
Figure 5: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean, Kendall’s $\tau$ for top-level demographic groups on three violation types when using experts to obtain binary labels $U$. All confidence intervals are within ±$\pm 0.01$.
@@ 5.3 Results by violation types
We first look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent. Figure 8 presents $Precision(S)$ and $Recall(S)$ for the crowd rater population on the three violation types when using experts to obtain binary labels $U$. Furthermore, figure 4 gives monotonic precision area, weighted recall area, their harmonic mean, and Kendall’s $\tau$ for the three violation types. The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since the severity of bias is harder to judge objectively. We also compare the scoring patterns of top-level demographic groups on each violation type. Figure 5 gives the metrics for top-level demographic groups on three violation types, using experts to obtain binary labels $U$. We see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations.

### v4

@@ 4.1 Metrics for the two properties
We formulate metrics for the two properties based on the standard concepts of precision and recall. Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, we take $Precision(S)$ to denote the precision when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe. Similarly, we take $Recall(S)$ to denote the recall when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe. The decision to compute precisions and recalls exactly at $S$ rather than $\geq S$ is a crucial one to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$. We define the following metrics to quantify the strength of the core inequalities for the two properties:
Monotonic Precision Area for stochastic ordering. We note that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$. We take the area under the curve defined by
$Y_{so}(s)=\sum_{i=0}^{s-1}\{Precision(s)-\max_{0\leq j\leq i}Precision(j)\}$ (3)
@@
at $s=0,1,2,3,\dots,K$, to directly and non-parametrically quantify increases in $P(U=1|S=s)$ while penalizing any violations of monotonicity at each score. Since $0\leq Precision(\cdot)\leq 1$, it can be shown via linear programming that the maximum possible area is given by $\left\lceil\frac{K+1}{2}\right\rceil*\left\lfloor\frac{K+1}{2}\right\rfloor$. To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold separating $U=1$ (unsafe) and $U=0$ (safe). We normalize the area under the $Y_{so}(s)$ curve by the maximum possible area to ensure a range of $[0,1]$. Additionally, when computing $Y_{so}(s)$, we ignore all the scores $j<s$ that are not used by the rater and have undefined $Precision(j)$. Similarly, if score $s$ is not used by the rater, we take $Y_{so}(s)=0$.
Weighted Recall Area for discrimination. We note that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe. Since $\sum Recall(\cdot)=1$, without needing normalization, we take the area under the curve defined by
$Y_{d}(s)=P(S<s|U=0)*Recall(s)$ (4)
@@
at $s=0,1,2,3,\dots,K$. A high area means high $P(S<s|U=0)$ and $P(S=s|U=1)$, which directly implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property. Essentially, $Y_{d}(s)$ is the recall of unsafe items at score $s$, i.e., $Recall(s)$, weighted by the proportion of safe items correctly discriminated by scores $<s$. It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to unsafe items while assigning scores $S<s$ to safe items.
Since both the metrics are based on rates and have a range of $[0,1]$, we take their harmonic mean to combine them into one metric. We have four cases to consider here. When a rater $j$ exhibits high MPA and high WRA, their $V^{j}$ is aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\approx V$, and they are responsive to both granular variations as well as broader distinctions in $V$. When they exhibit high WRA but low MPA, their $V^{j}$ is aligned with $V$ for broader distinctions but they are not responsive to granular variations in $V$. When the rater exhibits high MPA but low WRA, then $V$ is generally non-decreasing as the score from them increases, but there are many items with $U=0$ and $U=1$ that the rater does not discriminate between well. Finally, when the rater exhibits low MPA and low WRA, their $V^{j}$ is not aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\neq V$. We note that both the metrics are provably statistically consistent. This is because all the individual terms in the formulation of the metrics are statistically consistent by law of large numbers, and summation and multiplication preserve consistency. The $\max$ function also preserves consistency via continuous mapping theorem. Appendix B provides an implementation of our metrics in code.
@@ 4.2 Contrasting with traditional metrics
There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity. For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic. We highlight some limitations of these approaches below:
@@
  • Lack of focus on behaviors per score: correlation metrics capture general monotonic relationships between scores and underlying severity, if the latter were accessible, but they do not penalize raters who assign higher scores without meaningful increases in corresponding severity at the scores. Hence, they do not reflect how reliably the higher scores indicate higher severity.
  • Fragility to insignificant variations: traditional metrics focus solely on the ordering of scores relative to severity (ordinal relationships) but not the magnitude of differences between them (cardinal relationships). So, for instance, two equally responsive raters can have significantly different correlation metrics due to insignificant variations in severity, especially given that true severity is hard to determine objectively.
The way we define and quantify responsiveness addresses these limitations, while offering simplicity and the ability to get insights into how raters utilize different scores on the Likert scale. Nevertheless, as defined in our data model, different factors may influence how safety violations are perceived, regardless of the true underlying severity. While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity. In appendix C, we conduct extensive simulations to compare our proposed metrics against traditional metrics under different scoring patterns.
5 Evaluating and comparing responsiveness
To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to a publicly available pluralistic dataset of safety feedback on AI-generated content. Such safety alignment datasets with granular pluralistic feedback are not readily available.
@@ 5.1 Dataset description
We use the dataset of Rastogi et al. (2024), which contains safety feedback from diverse crowd raters for AI-generated images. Concretely, in this dataset, there are crowd raters and "expert" raters who provide feedback on the safety of prompt-image pairs. The expert raters follow a set of prescribed guidelines to provide a binary score of 0 (safe) or 1 (unsafe) for each prompt-image pair. They correspond to trained raters in our framework. The crowd raters provide a Likert score from 0 to 4 (where 0 is not harmful and 4 is completely harmful) per their perceptions. They were recruited based on their demographics. Rastogi et al. identified three demographic axes - gender, ethnicity and age. The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group. The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e. their ethnicity, age, and gender. Rastogi et al. distinguish top-level demographic groups, e.g., East-Asian, from trisectional demographic groups, e.g. Black–GenZ–Man. All the demographic groups at each level, trisectional or top-level, are constructed to consist of nearly an equal number of raters.
The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair. For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score. Taking the plurality vote preserves the ordinal nature of the scale. We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew. The distribution of scores provided by the crowd raters is shown in appendix D along with group cohesion metrics for each of the top-level demographic groups. Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
Figure 1 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference, and the grouping is by (a) ethnicity, (b) age, and (c) gender.
(a) Ethnicity
(b) Age
(c) Gender
Figure 1: $Precision(S)$ and $Recall(S)$ curves for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
@@ 5.2 Results by trisectional demographic groups
We now evaluate and compare the responsiveness of different demographic groups of crowd raters, both at the trisectional demographic level (e.g., Latinx–GenZ–Man) as well as the top demographic level (e.g., Latinx, Man, etc.). Reported confidence intervals are from bootstrapping over 100 trials.
@@
Figure 2: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
@@
Figure 2 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters. Figure 1(a) further validates the same as we note for instance that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(S)$ when plurality score $S$ goes from 1 to 3. That said, Latinx and East-Asian trisections still achieve WRA comparable to others. This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others. We note that Kendall’s $\tau$ and AUROC track WRA closely as both of them capture aspects of concordance. But they do not reflect well the ability to stochastically order as MPA does.
Figure 3 shows metrics for the trisectional demographic groups when the binary reference $U$ is obtained from crowd raters themselves, excluding the group being evaluated. We binarize the scores of crowd raters using boundaries $k\in[1,4]$ and report the metrics aggregated over all the boundaries via micro-averaging to ensure robustness. Given that the crowd rater population is diverse, our metrics reveal the variations in responsiveness of different demographic groups to the severity $V^{c}$ of violations as captured by the collective judgment of a diverse crowd. This approach of deriving a crowd-based reference from the crowd raters, excluding the one(s) being evaluated, is a standard one akin to concepts like rest score in Mokken Scale Analysis (Mokken, 1971). Here, we also provide metrics from Item Response Theory (IRT) and Mokken Scale Analysis (MSA) with the unidimensional latent trait being the true severity of prompt-image pairs. They are now applicable because now we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, Mokken h, and irt discrimination $\alpha$ for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from crowd raters, excluding the group being evaluated. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
MPA, WRA, and the traditional metrics continue to exhibit the same behaviors as before. Again, IRT and MSA only provide insights into raters’ ability to discriminate between levels of the latent trait based on the discrimination $\alpha$ or H values, but not the ability to stochastically order. Unlike before, we note that now the trisectional groups comprising Black ethnicity, aka Black trisections, consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05). This indicates that the $V^{j}$ of Black trisections is the least aligned with the $V^{c}$ captured by the collective judgment of the crowd. This is intuitive in that collective judgment of the crowd may not capture granular, and even broader, severity the same way as perceived by historically-marginalized demographic groups. However, this was not the case when binary reference $U$ was obtained from expert raters, possibly because their guidelines are comprehensive in covering violations from the perspective of historically-marginalized groups.
@@ 5.3 Results by violation types
We now look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent, when binary reference $U$ is obtained from expert raters. Figure 4 presents $Precision(S)$ and $Recall(S)$ curves for the three violation types and figure 4 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types. The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively. So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines. Appendix E further provides metrics for top-level demographic groups on violation types, and appendix F contains some qualitative examples.
Figure 4: (a) $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population. All confidence intervals are within ±$\pm 0.01$.
6 Recommendations and limitations
Gathering granular feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Rastogi et al., 2024; Kirk et al., 2024). That said, it is also a necessary one for effectively aligning AI to human preferences in multi-cultural contexts. Granular safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process:
  1. 1.
    To validate and improve guideline-driven safety alignment. Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (top-level, trisectional, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low. Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: identify raters who exhibit the highest responsiveness to severity from each and every group (top-level, trisectional, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
  2. 2.
    To directly perform safety alignment. Here, our findings show that naively optimizing for the collective judgment of the crowd may not effectively reflect the perceptions of all the groups. Firstly, our metrics can identify the groups whose granular or broader perceptions of severity may not be captured well in the collective judgment of a diverse crowd, i.e., low MPA or low WRA or both, be it collective judgment that is aggregated naively or otherwise (e.g., social choice). Further, our metrics can identify groups who show similar responsiveness to severity, which can potentially reduce the need to gather feedback from all such groups, and hence, reduce the costs.
We note that our metrics are descriptive and non-parametric in nature. So, while they enable a deeper interpretation of nuanced differences in observed trends, they cannot directly estimate the causal parameters behind the trends. Such estimations would require bigger datasets and more complex structural modeling, both of which we leave to future works.